OpenStudy (anonymous):

Write the equation of a line in point slope form that has a slope of 4 3 − and passes through the point (4, 5).

OpenStudy (anonymous):

-3/4*****

jigglypuff314 (jigglypuff314):

point slope form: y - y1 = m(x - x1) when m = slope and given the point (x1, y1) so plug in m = -3/4 and y1 = 5 and x1 = 4